OpenStudy (anonymous):

Verify the identity. tan^5x = tan^3xsec^2x -tan^3x

OpenStudy (anonymous):

\[tan^5x=tan^3x*tan^2x\] \[sin^2x+cos^2x=1\] \[tan^2x+1=sec^2x\] \[tan^2x=sec^2x-1\]

OpenStudy (anonymous):

\[tan^3x(sec^2x-1\]

OpenStudy (anonymous):

I got it! :) Thank you!